OpenStudy (anonymous):

show that the lines given by 2x^2 – 3xy – 2y^2 – 3x – 9y – 9 = 0 , 2x^2 – 3xy – 2y^2 = 0 are along the sides of a square ?

OpenStudy (anonymous):

factorise 1st equation we get (x-2y-3)(2x+y+3)=0 and 2nd give (x-2y)(2x+y)=0. so we get 4 equation . from here x-2y-3=0, x-2y=0 and 2x+y+3=0,2x+y=0 are parallel.